The sum of two numbers is 78. If three times the small number is subtracted from the larger number, the result is 14. Find the two numbers.
Answer by JackWestland(1)

The sum of two numbers is 78. If three times the small number is subtracted from the larger number, the result is 14. Find the two numbers.
X+Y=78 , Equation 1
X-3Y=14 , Equation 2
Equation 1-Equation 2
(X+Y=78) - (X-3Y=14)
X+Y)-(X-3Y) = 78)-(14)
X+Y-X+3Y = 78-14
4Y = 64
Y= 16
X+Y=78
78-Y=X
78-16=X
X=62
X=62 Y=16